### v2

@@ Abstract
Understanding people’s preferences and needs is crucial for urban planning decisions, yet current approaches often combine them from multi-cultural and multi-city populations, obscuring important demographic differences and risking amplifying biases. We conducted a large-scale urban visual perception survey of streetscapes worldwide using street view imagery, examining how demographics – including gender, age, income, education, race and ethnicity, and, for the first time, personality traits – shape perceptions among 1,000 participants, with balanced demographics, from five countries and 45 nationalities. This dataset, introduced as Street Perception Evaluation Considering Socioeconomics (SPECS), exhibits statistically significant differences in perception scores in six traditionally used indicators (safe, lively, wealthy, beautiful, boring, and depressing) and four new ones we propose (live nearby, walk, cycle, green) among demographics and personalities. We revealed that location-based sentiments are carried over in people’s preferences when comparing urban streetscapes with other cities. Further, we compared the perception scores based on where participants and streetscapes are from. We found that an off-the-shelf machine learning model trained on an existing global perception dataset tends to overestimate positive indicators and underestimate negative ones compared to human responses, suggesting that targeted intervention should consider locals’ perception. Our study aspires to rectify the myopic treatment of street perception, which rarely considers demographics or personality traits.
@@ 2.1 Demographics and personalities as moderators of perception
Figure 2: Statistical difference of perception scores by demographics. Perception Q scores are calculated from ratings by participants in each subgroup for all (without location grouping) and each location. Locations with no significant differences in any indicator and demographic groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per subgroup) are not shown. Significance thresholds *$p<0.05$.
We evaluated whether the mean SVI perception Q scores significantly differ across participants grouped by demographic characteristics and personality traits. We found differences between demographic and personalities subgroups but they differ by locations, i.e., where participants are from, and perceptual indicators (Figure 2). The highest number of significant differences happened in populations grouped by gender, and across all five demographics (gender, age group, annual household income bracket, education level, and race and ethnicity) the traditional indicator safe had the most intra-country differences, followed by our new indicator cycle. No significant differences were found in negative indicators and only one significant difference was found in the live nearby and lively indicator in gender and race and ethnicity subgroups, respectively, in the USA. In terms of countries, we found differences in participants from San Francisco and Santa Clara (USA) in four out of the five demographic characteristics and no differences in any demographic group of participants from Santiago (Chile) or Singapore. We noted gender differences under the live nearby, and cycle indicators for specific countries (USA and Nigeria, respectively), and across all locations (e.g., Q scores calculated without grouping by countries) under the safe indicator (top plot in Figure 2).
@@
While overall we found fewer differences within demographic interaction subgroups, i.e., a total of four compared to eight in single demographic groups (Figure 2), among all the combinations of ‘gender $\times$ age group $\times$ AHI’ subgroups, half the differences happened within-country and half across all locations (Figure 9). Differences formed by all three demographics combined were found for the lively indicator between “Female x 30-39 x Bottom” and “Male x 21-29 x Middle” subgroups only when all participants’ responses were combined, although under a small sample size ($n\geq 5$, top plot in Figure 9). Within ‘gender $\times$ age group’ subgroups perception differences for the walk indicator between genders of the “Above 50” age group in the USA also under a smaller sample size ($n\geq 5$, second plot from the top Figure 9). Then, differences for the safe indicator between genders of the “Middle” AHI group were found in Netherlands under a larger sample size. Finally, perceptual differences between age groups “21-29” and “30-39” of the “Bottom” AHI bracket were found for the lively indicator across all locations. Both perceptual differences for safe and lively were found with higher sample size ($n\geq 20$, last two bottom plots Figure 9, respectively).
Figure 3: Statistical difference of perception scores by personality traits. Perception Q scores are calculated from ratings by participants in each subgroup for all (without location grouping) and each location. Locations with no significant differences in any indicator and personality groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per subgroup) are not shown. Significance threshold *$p<0.05$.
@@ 2.4 Perceptual variation from residents vs non-residents
As the last comparison of perception Q scores, we expanded on the second analysis on multiple-city SVI and single-city participants and looked into how people living in the same, or neighboring, city perceive other cities relative to their own. We analyzed the average change in the perception scores, scaled to z-scores, of participants who rated the SVI of their own city compared to the SVI of the other four cities and found that some populations perceived their city mostly better (higher positive indicators scores) and others perceived their city overall worse (lower positive indicator and higher negative indicator scores) (bottom plot in Figure 4). In this figure, the differences shown on the slopes are based on z-scores and allow us to compare how different a response is when compared to all responses, i.e., standard deviation units (SDs). This format allows us to avoid the participants’ bias that could appear when using absolute perception Q scores.
@@ 3 Discussion
Across all cities chosen, the statistical analysis revealed that differences in global urban visual perception are explained by demographics and personalities (Figure 2). While existing work found safety perceptions, in specific locations, are explained by gender [34, 35, 15, 14, 16], our findings revealed a potential aggregated difference when considering all five locations together. Similar to work done in [11], we did not find statistically significant differences between age groups for safety perceptions, although the age group representation in said work was less granular: below or above 50 years old. Another work also did not find significant differences in perception scores by age or gender [9], although similarly to [11], age groups were analyzed as below or above a given age, i.e., 28 years old in said study. Unlike work in [34], we did not find significant differences in the boring indicator for gender, or any, demographic subgroups. Nevertheless, our findings show that in other indicators, both traditional and new, there are significant differences for all five demographic characteristics in different countries. More importantly, these significant differences are found in cities from continents often overlooked by the research community, e.g., Africa [36]. Furthermore, although nested demographic groups presented limitations in sample size, we found significant differences with reasonable sample sizes ($n\geq 20$) in groups that included AHI (Figure 9).
Our findings on personality traits follow this same trend (Figure 3). With a similar amount of differences compared to those observed across single demographic subgroups, Abuja (Nigeria), Singapore, and San Francisco and Santa Clara (USA) exhibit differences among residents with distinct personality traits, e.g., personality scores in either Q1 or Q3 of their cities. Personality traits have been linked with subjective well-being, for example, conscientiousness was found to be a good predictor of life satisfaction [37]. Personality subgroups based on this, and the extraversion trait, led to the most significant differences including when all participants’ ratings were combined (middle plot in Figure 3). These results open the opportunity to quantify how urban streetscapes are perceived by residents of different levels of well-being. Despite the remaining three personality traits showing only one significant association, our overall findings reinforce ample literature correlating personality and landscape preference and the relationship of personality and aesthetic judgment [38]. These links are crucial when designing tailored urban solutions as most of the significant differences over demographic and personality profiles were found within countries, 15 differences compared to four differences across all locations.
@@
Despite the demographic diversity and geographic location of participants, most of our proposed indicators (live nearby, cycle, and walk) capture, predominantly linearly, the trends of the existing six indicators. These new dimensions are able to expand perception studies with indicators reflecting a user’s willingness to be associated with an urban scene (e.g., live, walk, or cycle in the surroundings). The green indicator introduces the opportunity to expand greenery assessment from traditional objective measurements such as pixel-ratio metrics, e.g., Green View Index (GVI), to subjective context-aware human perception data. While the relationships found with other perceptual indicators were more complex and nonlinear, we found around half of the statistically significant differences in perception scores between demographic and personality profiles (6 out of 13) under these four new indicators. These highlight the potential of these indicators as input for targeted interventions. As different demographic subgroups in diverse location perceive them differently; e.g., different live nearby and cycle preferences between gender groups in San Francisco and Santa Clara (USA), and Abuja (Nigeria), respectively (Figure 2); different green preferences between high and low agreeableness groups in Singapore; and different walk preferences in genders of a given age group in San Francisco and Santa Clara (USA) (Figure 9). These indicators offer a clearer human subjective perspective that, when complemented with objective measurements, leads to improved urban solutions and assessments [42, 11, 33]. We propose that researchers consider including these indicators where possible and when relevant.
@@
This work contributes to the few studies in multi-city perception, emphasizes the importance of human-centric data collection, and releases a global and diverse perception dataset. Methodologically, we introduced a global and, participants-wise, demographically balanced urban visual perception called SPECS and we thoroughly analyzed perception scores by demographic and personality subgroups. Moreover, we quantified and statistically verified, for the first time, the magnitude bias of off-the-shelf machine learning models in perceptual predictions despite being trained on a global dataset like PP2, we isolated and analyzed the effect of how people are influenced by preconceived conceptions of urban contexts, and introduced and validated new indicators relevant to the six commonly used. Our results draws us closer to understanding in what contexts, e.g., locations and perceptual indicators, urban visual perception is most affected by demographics and personality. While we are not able to pinpoint a specific number or percentage of how much people’s profiles drive perception responses, we found that as such it can lead to significant differences depending on the profiles and locations. These insights on profiles and location subgroups can be extended to other urban studies, such as building exterior evaluation for urban development and strategies [43]. Street-level image-based indicators closely correlate with a city’s social structure and economic development [44]. Thus, as SVI captures the human-level perspective, more closely aligned with people’s experiences in cities, studying and understanding urban streetscapes allows researchers and stakeholders to take informed actions [45]. While PP2 remains an impressive effort valuable for urban science, our analysis reveals its limitations in localized contexts, suggesting that generalized or one-size-fits-all models may produce distorted results. We recommend that researchers fine-tune their perception prediction models with local data, even though this may require additional data collection efforts. This approach is likely to produce more accurate and better performing models [11, 16].
@@
Secondly, we acknowledge the limited number of cities and countries chosen as well as the respective number of images. Our approach of one city per country per continent limits the generalizability of our findings to the broader urban landscape worldwide. To mitigate these limitations, we employed stratified sampling of both urban environments (image selection) and participant demographics, ensuring proportional representation of local socioeconomic contexts within each selected city. Additionally, to ensure statistical power across all ten different indicators for demographic and personality subgroups’ perception ratings analyses, we constrained the number of images per city to 80, to collect enough participant ratings per image and per subgroup to support robust comparisons across multiple demographic and personality dimensions. Rather than claiming global or country-level representativeness and insights, our study focuses on determining whether perceptual indicators vary systematically across demographic groups in diverse urban environments. The reduced number of images per city, taken primarily from city centers, might introduce location biases that could affect the interpretation of our findings. These location biases may include overrepresentation of tourist areas, commercial districts, or architecturally distinctive features that may not reflect typical residential or suburban environments within each city, e.g., very unique urban forms that people associate with specific cities, such as Amsterdam’s canals. While our stratified sampling approach helps ensure demographic representativeness within each city, it cannot fully address the geographical limitations of our urban sampling strategy. Nevertheless, our contributions rely on thorough comparisons of perceptions across different demographic subgroups (top and middle plots in Figure 2), with analyses that control for location when possible (bottom plot in Figure 4). This methodological approach allows us to isolate demographic effects on urban perception across multiple diverse urban contexts, i.e., five cities in five different continents compared to perception studies routinely using one or two cities [5, 31, 32, 33].
@@
Future research should prioritize analyzing interactions between multiple demographic factors, e.g., older adults with different walk abilities [46], and their combined effects on various perceptual indicators. As our study nested participant subgroups by demographics, the resulting smaller participant pools yielded fewer ratings per perceptual indicator (e.g., safety perception scores based solely on older women’s ratings). To address this limitation, subsequent studies could either increase participant numbers or ratings per image. The latter could be achieved by reducing the number of images per city, though potentially compromising urban landscape diversity, or by focusing on fewer perceptual indicators, such as exclusively studying safety perception. Another alternative is to leverage scalable perception rating methods. Human-machine frameworks [47] allows for rapid and cost-effective perception assessments. Building on our findings on the demographic and location-specific perceptual differences, these methods could be more appropriately deployed to quickly evaluate entire cities.
@@ 4.2 Perception survey
The survey was reviewed and approved by the Institutional Review Board (IRB) of our university, and the participants were compensated financially. It was deployed on an online platform and distributed by a third-party local market research vendor. Using a single data collection vendor with global presence allowed us to recruit 200 participants from each of the five countries mentioned in the above subsection and ensure validity of data and demographic balance, and guarantee consistency in recruiting and data collection. Considering the number of images, this number of participants was chosen primarily to gather enough demographic subgroups for downstream analyses of their survey responses. The complete survey consisted of two different sections; the first section captured the demographics in 11 questions, and the second consisted of 50 unique pairwise comparisons of the 400 images across 10 subjective indicators. Image 2 in the bottom panel of Figure 1 shows the different demographic moderators and human perception rating indicators.
@@ 4.4 Data analysis
We divided the analysis into five parts, as noted in the bottom panel of Figure 1 step 4. First, we looked at demographic factors and personality traits as moderators in urban visual perception. For each country, we created subgroups based on the available answers for each demographic question: gender, age group, annual household income (AHI), education level, and race and ethnicity. While we tried to keep the original groups on each demographic question, the granularity in some answers presented challenges. The available groups in AHI are as few as five groups and as many as 12 groups and are different in each country, which would also make it difficult to compare answers between countries. Thus, we decided to remap existing answers into Lower, Middle, and Upper groups (Table 2). We also remapped some of the answer from the education level question into a common single one. Answers such as “Masters degree” and “Doctorate or professional degree” were remapped into “Postgraduate degree”.
With this variable remapping, we also explored demographics interactions where possible. Using all five demographic factors (i.e., gender, age, income, education, race and ethnicity) to create nested subgroup reduces the available pairwise comparisons to a maximum of one pairwise comparison per image across all indicators and locations. Thus, we used only the combinations of ‘gender $\times$ age group $\times$ AHI’ as nested subgroups as these demographic factors are commonly explored in urban perception studies [13, 14, 15, 71] and among our demographics factors, they have fewer categories; two, four, and three, respectively. Nested subgroups based on these demographic factors have an average of six pairwise comparisons per image across all indicators and locations, above the threshold of four pairwise comparisons.
When analyzing each demographic factor in isolation and in interactions, i.e., computing Q scores based on a demographic subgroup, we did not consider ratings from participants who answered “Prefer not to answer”(39 out of 1,000 participants in AHI), “Less than Secondary/High school” and “Other” (11 out of 1,000 participants in education level), and “A race/ethnicity not listed here” (69 out of 1,000 participants in the race and ethnicity question) due to their small sample size. On average, these excluded responses would have contributed only one additional rating to the Q score computation of each perceptual indicator for between 5 and 76 images per location. All subsequent analyses used all participants’ ratings.
@@
We created these demographic and personality subgroups for each country, and a set of subgroups by combining all countries too. While this nested subgrouping allowed for more detailed analysis, by reducing the number of available participants in each group, the number of pairwise comparisons per image per indicator was also reduced. Thus, after computing the Q scores in each nested subgroup, we retained all images with at least four pairwise comparisons, as this was the reported average comparison in PP2 [6]. Since each demographic subgroup has a different number of samples, i.e., rated images with at least four pairwise comparison by participants with the same demographic characteristics, we prioritized a higher minimum number of samples, i.e., n $\geq 20$, where applicable but reported the minimum used in each demographic subgroup.
To quantify the influence of demographic and personality attributes, we performed statistical tests on the Q scores for all SVIs rated by participants in their respective subgroups. Given the difference in sample sizes, i.e., the number of rated SVI by available participants in the nested subgroups, and normal distribution but unequal variance of Q scores, we chose Welch’s ANOVA as the statistical method to compare whether the subgroups’ perception scores means are significantly different. For demographic attributes with more than two groups, we performed the Games-Howell post-hoc test to identify the significant differences, if any, between each pair of groups’ means.
@@
Annotated and labeled SVI data are obtained from the NUS Global Streetscapes dataset (https://huggingface.co/datasets/NUS-UAL/global-streetscapes). Raw images were obtained from the crowdsourced platforms Mapillary and KartaView. We share openly our dataset SPECS (Street Perception Evaluation Considering Socioeconomics), consisting of survey responses and participants’ demographic data at [link will be added for publication].
@@
The step-by-step process and code for all analyses are available in the public repository: [link will be added for publication].

### v3

@@ Abstract
Understanding people’s preferences is crucial for urban planning, yet current approaches often combine responses from multi-cultural populations, obscuring demographic differences and risking amplifying biases. We conducted a large-scale urban visual perception survey of streetscapes worldwide using street view imagery, examining how demographics – including gender, age, income, education, race and ethnicity, and, for the first time, personality traits – shape perceptions among 1,000 participants with balanced demographics from five countries and 45 nationalities. This dataset, Street Perception Evaluation Considering Socioeconomics (SPECS), reveals demographic- and personality-based differences across six traditional indicators (safe, lively, wealthy, beautiful, boring, depressing) and four new ones (live nearby, walk, cycle, green). Location-based sentiments further shape these preferences. Machine learning models trained on existing global datasets tend to overestimate positive indicators and underestimate negative ones compared to human responses, underscoring the need for local context. Our study aspires to rectify the myopic treatment of street perception, which rarely considers demographics or personality traits.
@@ 2.1 Demographics and personalities as moderators of perception
Figure 2: Statistical difference of perception scores by demographics. Perception Q scores are calculated from ratings by participants in each group for all (without location grouping) and each location. Locations with no significant differences in any indicator and demographic groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per group) are not shown. Significance thresholds *$p<0.05$.
We evaluated whether the mean SVI perception Q scores significantly differ across participants grouped by demographic characteristics and personality traits. We found differences between demographic and personalities groups but they differ by locations, i.e., where participants are from, and perceptual indicators (Figure 2). The highest number of significant differences happened in populations grouped by gender, and across all five demographics (gender, age group, annual household income bracket, education level, and race and ethnicity) the traditional indicator safe had the most intra-country differences, followed by our new indicator cycle. No significant differences were found in negative indicators and only one significant difference was found in the live nearby and lively indicator in gender and race and ethnicity groups, respectively, in the USA. In terms of countries, we found differences in participants from San Francisco and Santa Clara (USA) in four out of the five demographic characteristics and no differences in any demographic group of participants from Santiago (Chile) or Singapore. We noted gender differences under the live nearby, and cycle indicators for specific countries (USA and Nigeria, respectively), and across all locations (e.g., Q scores calculated without grouping by countries) under the safe indicator (top plot in Figure 2).
@@
While overall we found fewer differences within demographic interaction nested groups, i.e., a total of four compared to eight in single demographic groups (Figure 2), among all the combinations of ‘gender $\times$ age group $\times$ AHI’ nested groups, half the differences happened within-country and half across all locations (Figure 9). Differences formed by all three demographics combined were found for the lively indicator between “Female x 30-39 x Bottom” and “Male x 21-29 x Middle” nested groups only when all participants’ responses were combined, although under a small sample size ($n\geq 5$, top plot in Figure 9). Within ‘gender $\times$ age group’ nested groups perception differences for the walk indicator between genders of the “Above 50” age group in the USA also under a smaller sample size ($n\geq 5$, second plot from the top Figure 9). Then, differences for the safe indicator between genders of the “Middle” AHI group were found in Netherlands under a larger sample size. Finally, perceptual differences between age groups “21-29” and “30-39” of the “Bottom” AHI bracket were found for the lively indicator across all locations. Both perceptual differences for safe and lively were found with higher sample size ($n\geq 20$, last two bottom plots Figure 9, respectively).
Figure 3: Statistical difference of perception scores by personality traits. Perception Q scores are calculated from ratings by participants in each group for all (without location grouping) and each location. Locations with no significant differences in any indicator and personality groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per group) are not shown. Significance threshold *$p<0.05$.
@@ 2.4 Perceptual variation from residents vs non-residents
As the last comparison of perception Q scores, we expanded on the second analysis on multiple-city SVI and single-city participants and looked into how people living in the same, or neighboring, city perceive other cities relative to their own. We analyzed the average change in the perception scores, scaled to z-scores within participants’ locations to debias responses from cultural and location-based influences, of participants who rated the SVI of their own city compared to the SVI of the other four cities. We found that some populations perceived their city mostly better (higher positive indicators scores) and others perceived their city overall worse (lower positive indicator and higher negative indicator scores) (bottom plot in Figure 4). In this figure, the differences shown on the slopes are based on z-scores and allow us to compare how different a response is when compared to all responses, i.e., standard deviation units (SDs). This format allows us to avoid the participants’ bias that could appear when using absolute perception Q scores.
@@ 3 Discussion
Across all cities chosen, the statistical analysis revealed that differences in global urban visual perception are explained by demographics and personalities (Figure 2). While existing work found safety perceptions, in specific locations, are explained by gender [34, 35, 15, 14, 16], our findings revealed a potential aggregated difference when considering all five locations together. Similar to work done in [11], we did not find statistically significant differences between age groups for safety perceptions, although the age group representation in said work was less granular: below or above 50 years old. Another work also did not find significant differences in perception scores by age or gender [9], although similarly to [11], age groups were analyzed as below or above a given age, i.e., 28 years old in said study. Unlike work in [34], we did not find significant differences in the boring indicator for gender, or any, demographic groups. Nevertheless, our findings show that in other indicators, both traditional and new, there are significant differences for all five demographic characteristics in different countries. More importantly, these significant differences are found in cities from continents often overlooked by the research community, e.g., Africa [36]. Furthermore, although nested demographic groups presented limitations in sample size, we found significant differences with reasonable sample sizes ($n\geq 20$) in groups that included AHI (Figure 9).
Our findings on personality traits follow this same trend (Figure 3). With a similar amount of differences compared to those observed across single demographic groups, Abuja (Nigeria), Singapore, and San Francisco and Santa Clara (USA) exhibit differences among residents with distinct personality traits, e.g., personality scores in either Q1 or Q3 of their cities. Personality traits have been linked with subjective well-being, for example, conscientiousness was found to be a good predictor of life satisfaction [37]. Personality groups based on this, and the extraversion trait, led to the most significant differences including when all participants’ ratings were combined (middle plot in Figure 3). These results open the opportunity to quantify how urban streetscapes are perceived by residents of different levels of well-being. Despite the remaining three personality traits showing only one significant association, our overall findings reinforce ample literature correlating personality and landscape preference and the relationship of personality and aesthetic judgment [38]. These links are crucial when designing tailored urban solutions as most of the significant differences over demographic and personality profiles were found within countries, 15 differences compared to four differences across all locations.
@@
Despite the demographic diversity and geographic location of participants, most of our proposed indicators (live nearby, cycle, and walk) capture, predominantly linearly, the trends of the existing six indicators. These new dimensions are able to expand perception studies with indicators reflecting a user’s willingness to be associated with an urban scene (e.g., live, walk, or cycle in the surroundings). The green indicator introduces the opportunity to expand greenery assessment from traditional objective measurements such as pixel-ratio metrics, e.g., Green View Index (GVI), to subjective context-aware human perception data. While the relationships found with other perceptual indicators were more complex and nonlinear, we found around half of the statistically significant differences in perception scores between demographic and personality profiles (6 out of 13) under these four new indicators. These highlight the potential of these indicators as input for targeted interventions. As different demographic groups in diverse location perceive them differently; e.g., different live nearby and cycle preferences between gender groups in San Francisco and Santa Clara (USA), and Abuja (Nigeria), respectively (Figure 2); different green preferences between high and low agreeableness groups in Singapore; and different walk preferences in genders of a given age group in San Francisco and Santa Clara (USA) (Figure 9). These indicators offer a clearer human subjective perspective that, when complemented with objective measurements, leads to improved urban solutions and assessments [42, 11, 33]. We propose that researchers consider including these indicators where possible and when relevant.
@@
This work contributes to the few studies in multi-city perception, emphasizes the importance of human-centric data collection, and releases a global and diverse perception dataset. Methodologically, we introduced a global and, participants-wise, demographically balanced urban visual perception called SPECS and we thoroughly analyzed perception scores by demographic and personality groups. Moreover, we quantified and statistically verified, for the first time, the magnitude bias of off-the-shelf machine learning models in perceptual predictions despite being trained on a global dataset like PP2, we isolated and analyzed the effect of how people are influenced by preconceived conceptions of urban contexts, and introduced and validated new indicators relevant to the six commonly used. Our results draws us closer to understanding in what contexts, e.g., locations and perceptual indicators, urban visual perception is most affected by demographics and personality. While we are not able to pinpoint a specific number or percentage of how much people’s profiles drive perception responses, we found that as such it can lead to significant differences depending on the profiles and locations. These insights on profiles and location groups can be extended to other urban studies, such as building exterior evaluation for urban development and strategies [43]. Street-level image-based indicators closely correlate with a city’s social structure and economic development [44]. Thus, as SVI captures the human-level perspective, more closely aligned with people’s experiences in cities, studying and understanding urban streetscapes allows researchers and stakeholders to take informed actions [45]. While PP2 remains an impressive effort valuable for urban science, our analysis reveals its limitations in localized contexts, suggesting that generalized or one-size-fits-all models may produce distorted results. We recommend that researchers fine-tune their perception prediction models with local data, even though this may require additional data collection efforts. This approach is likely to produce more accurate and better performing models [11, 16].
@@
Secondly, we acknowledge the limited number of cities and countries chosen as well as the respective number of images. Our approach of one city per country per continent limits the generalizability of our findings to the broader urban landscape worldwide. To mitigate these limitations, we employed stratified sampling of both urban environments (image selection) and participant demographics, ensuring proportional representation of local socioeconomic contexts within each selected city. Additionally, to ensure statistical power across all ten different indicators for demographic and personality groups’ perception ratings analyses, we constrained the number of images per city to 80, to collect enough participant ratings per image and per group to support robust comparisons across multiple demographic and personality dimensions. Rather than claiming global or country-level representativeness and insights, our study focuses on determining whether perceptual indicators vary systematically across demographic groups in diverse urban environments. The reduced number of images per city, taken primarily from city centers, might introduce location biases that could affect the interpretation of our findings. These location biases may include overrepresentation of tourist areas, commercial districts, or architecturally distinctive features that may not reflect typical residential or suburban environments within each city, e.g., very unique urban forms that people associate with specific cities, such as Amsterdam’s canals. While our stratified sampling approach helps ensure demographic representativeness within each city, it cannot fully address the geographical limitations of our urban sampling strategy. Nevertheless, our contributions rely on thorough comparisons of perceptions across different demographic groups (top and middle plots in Figure 2), with analyses that control for location when possible (bottom plot in Figure 4). This methodological approach allows us to isolate demographic effects on urban perception across multiple diverse urban contexts, i.e., five cities in five different continents compared to perception studies routinely using one or two cities [5, 31, 32, 33].
@@
Future research should prioritize analyzing interactions between multiple demographic factors, e.g., older adults with different walk abilities [46], and their combined effects on various perceptual indicators. As our study grouped participants by multiple demographics factors, the resulting smaller participant pools yielded fewer ratings per perceptual indicator (e.g., safety perception scores based solely on older women’s ratings). To address this limitation, subsequent studies could either increase participant numbers or ratings per image. The latter could be achieved by reducing the number of images per city, though potentially compromising urban landscape diversity, or by focusing on fewer perceptual indicators, such as exclusively studying safety perception. Another alternative is to leverage scalable perception rating methods. Human-machine frameworks [47] allows for rapid and cost-effective perception assessments. Building on our findings on the demographic and location-specific perceptual differences, these methods could be more appropriately deployed to quickly evaluate entire cities.
@@ 4.2 Perception survey
The survey was reviewed and approved by the Institutional Review Board (IRB) of our university, and the participants were compensated financially. It was deployed on an online platform and distributed by a third-party local market research vendor. Using a single data collection vendor with global presence allowed us to recruit 200 participants from each of the five countries mentioned in the above subsection and ensure validity of data and demographic balance, and guarantee consistency in recruiting and data collection. Considering the number of images, this number of participants was chosen primarily to gather enough demographic groups for downstream analyses of their survey responses. The complete survey consisted of two different sections; the first section captured the demographics in 11 questions, and the second consisted of 50 unique pairwise comparisons of the 400 images across 10 subjective indicators. Image 2 in the bottom panel of Figure 1 shows the different demographic moderators and human perception rating indicators.
@@ 4.4 Data analysis
We divided the analysis into five parts, as noted in the bottom panel of Figure 1 step 4. First, we looked at demographic factors and personality traits as moderators in urban visual perception. For each country, we created groups based on the available answers for each demographic question: gender, age group, annual household income (AHI), education level, and race and ethnicity. While we tried to keep the original groups on each demographic question, the granularity in some answers presented challenges. The available groups in AHI are as few as five groups and as many as 12 groups and are different in each country, which would also make it difficult to compare answers between countries. Thus, we decided to remap existing answers into Lower, Middle, and Upper groups (Table 2). We also remapped some of the answer from the education level question into a common single one. Answers such as “Masters degree” and “Doctorate or professional degree” were remapped into “Postgraduate degree”.
With this variable remapping, we also explored demographics interactions where possible. Using all five demographic factors (i.e., gender, age, income, education, race and ethnicity) to create nested groups greatly reduces the number of participants within each group. This reduction limits the available pairwise comparisons to a maximum of one pairwise comparison per image across all indicators and locations. Therefore, we used only the combinations of ‘gender $\times$ age group $\times$ AHI’ as nested groups. These demographic factors are commonly explored in urban perception studies [13, 14, 15, 71] and among our demographic factors, they have the fewest categories; two, four, and three, respectively; resulting in fewer nested groups. Nested groups based on these three demographic factors yield an average of six pairwise comparisons per image across all indicators and locations, exceeding our selected threshold of four pairwise comparisons used throughout our analyses.
When analyzing each demographic factor in isolation and in interactions, i.e., computing Q scores based on a demographic group, we did not consider ratings from participants who answered “Prefer not to answer”(39 out of 1,000 participants in AHI), “Less than Secondary/High school” and “Other” (11 out of 1,000 participants in education level), and “A race/ethnicity not listed here” (69 out of 1,000 participants in the race and ethnicity question) due to their small sample size. On average, these excluded responses would have contributed only one additional rating to the Q score computation of each perceptual indicator for between 5 and 76 images per location. All subsequent analyses used all participants’ ratings.
@@
We created these demographic and personality groups for each country, and a set of groups by combining all countries too. While this nested grouping allowed for more detailed analysis, by reducing the number of available participants in each group, the number of pairwise comparisons per image per indicator was also reduced. Thus, after computing the Q scores in each nested group, we retained all images with at least four pairwise comparisons, as this was the reported average comparison in PP2 [6]. Since each demographic group has a different number of samples, i.e., rated images with at least four pairwise comparison by participants with the same demographic characteristics, we prioritized a higher minimum number of samples, i.e., n $\geq 20$, where applicable but reported the minimum used in each demographic group.
To quantify the influence of demographic and personality attributes, we performed statistical tests on the Q scores for all SVIs rated by participants in their respective groups. Given the difference in sample sizes, i.e., the number of rated SVI by available participants in each group, and normal distribution but unequal variance of Q scores, we chose Welch’s ANOVA as the statistical method to compare whether the groups’ perception scores means are significantly different. For demographic attributes with more than two groups, we performed the Games-Howell post-hoc test to identify the significant differences, if any, between each pair of groups’ means.
@@
Annotated and labeled SVI data are obtained from the NUS Global Streetscapes dataset (https://huggingface.co/datasets/NUS-UAL/global-streetscapes). Raw images were obtained from the crowdsourced platforms Mapillary and KartaView. We share openly our dataset SPECS (Street Perception Evaluation Considering Socioeconomics), consisting of survey responses and participants’ demographic data at https://huggingface.co/datasets/matiasqr/specs.
@@
The step-by-step process and code for all analyses are available in the public repository: https://github.com/matqr/specs.